• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Boeglin

    Log entry time 00:01:58 on December 1,2007

    Entry number 217322

    keyword=Shift summary

    16:00 waiting for beam
    18:10 MCC ask for permission to put CW beam for Harp scans
    18:20 harp scan results (A) : 168 x 121 um, (B) 156 x 99 um
    18:58 goto beam left Fe for first production run
    19:00 compton rate unacceptable, call MCC. They are re-tuning
    20:25 MCC still working on adjusting the beam size
    20:45 Harp scan results: (A) 208 x 152, (B) 77 x 130
    21:12 Beam back , Compton looks good
    21:20 run 3072: beam spot too narrow. Increase raster=>compton rate up again (40 - 50k)
    22:05 Harp scan results : (A) 251x150 (B) 77x134
    22:07 DAQ crashed
    22:25 beam spot looks better. Increase raster by anoth factor 1.5
    22:35 raster good MCC setting : 6.3 x start
    start production again